(13) x1 = 4x1 - x2 + 3e2t x1(0) = - x2 = 2x1 + x2 + 2t X2(0) = 42

Differential Equations Question

Method of Elimination - Initial Value Problems.

Find the solution of the following IVPs.
Answer #1

Sol à = 42, - ng+ 3 e at ng= an, eng & ait and x = usi, -ig + beant * = 42-(022, ++4 ) +6edct * = 45,- (an, #3, +44, +3e28 +2and porro na = -; +421 + 300 € = -3c, eget – ao cart + 3e act+ axa de act + t uc, eget tuc eget - 12 et eget hogy -20 +3 eact